# LUMOS: Large User MOdel Series for User Behavior Prediction

### v1

@@ Abstract
User behavior prediction at scale remains a critical challenge for online B2C platforms. Traditional approaches rely heavily on task-specific models and domain-specific feature engineering. This is time-consuming, computationally expensive, and requires domain expertise and therefore not scalable. We present LUMOS (Large User MOdel Series), a transformer-based architecture that eliminates task-specific models and manual feature engineering by learning multiple tasks jointly using only raw user activity data. LUMOS introduces a novel cross-attention mechanism that conditions predictions on future known events (e.g., holidays, sales, etc.), enabling the model to predict complex behaviour patterns like ”how will upcoming holidays affect user engagement?” The architecture also employs multi-modal tokenization, combining user transactions, event context, and static user demographic attributes into rich representations processed through specialized embedding pathways.
Through extensive experiments on a production dataset spanning 275 billion user activity tokens from 250 million users, we demonstrate that LUMOS achieves superior performance compared to traditional task-specific models. Across 5 tasks with established baselines, we achieve an average improvement of 0.025 in ROC-AUC for binary classification tasks and 4.6% reduction in MAPE for regression tasks. Online A/B testing validates these improvements translate to measurable business impact with a 3.15% increase in Daily Active Users.
@@ 1 Introduction
We propose a shift in user behavior modelling inspired by the trends in NLP where the field moved from task-specific models and feature-based approaches (like bag of words and TF-IDF) to foundation models like BERT [1] and GPT [8] which are not just easier to train (in terms of pre-processing steps) but also achieve better performance when data scale is sufficient.
Figure 1: Paradigm shift from multiple specialized models to a single unified model.
In this spirit, we present LUMOS, a transformer-based architecture that can be trained to predict future user behaviour using raw sequence of past behaviour. It eliminates manual feature engineering by learning end-to-end representations from raw daily user activity data. Instead of predicting a single user behaviour attribute for a pre-specified time horizon, LUMOS predicts a sequence of user behaviour tokens for multiple future time steps. Each predicted token contains the full set of predicted user behaviours precluding the need for multiple specialized models. Experiments on our production dataset show that LUMOS achieves an average improvement of 0.025 in AUC for classification tasks and 4.6% reduction in MAPE for regression tasks compared to the baseline task-specific models.
LUMOS uses a cross-attention mechanism between historical user behavior and future event context to answer nuanced questions like ”Given that the IPL final is next Sunday, which days in the user’s history are most relevant for predicting their engagement?” Our experiments show that the model correctly attends not only to recent activity but also to specific historical events like previous year’s IPL finals or World Cup matches, learning these patterns entirely from data without explicit programming. This is especially crucial for event driven platforms like sports streaming platforms or ticketing services.
@@ 2 Background
User behavior modeling encompasses a range of tasks critical to online platforms: predicting churn, forecasting engagement levels, estimating customer lifetime value, and anticipating user actions [10]. These predictions typically span multiple time horizons (7-day, 28-day, annual) and serve different business functions from personalized marketing to risk management.
@@
Recent work has successfully applied transformers to sequential user behavior modeling. SASRec [4] pioneered the use of self-attention for sequential recommendation by modeling item sequences with left-to-right attention masks. BERT4Rec [13] adopted bidirectional encoder representations from BERT [1] for item prediction by masking and predicting random items in user interaction sequences. TiSASRec [7] extended this approach by incorporating explicit time interval information between consecutive interactions to capture temporal dynamics more effectively. However, these approaches primarily focus on single-modality item sequences and do not naturally accommodate heterogeneous data sources such as continuous numerical transaction features, external event context, or static user attributes. Furthermore, they employ decoder-only or masked prediction architectures that cannot leverage known future context—such as upcoming events or scheduled activities—that may influence user behavior.
Adapting transformers to comprehensive user behavior prediction thus presents several unique challenges beyond those addressed by existing sequential recommendation models:
1. Continuous vs. Discrete Temporal Patterns: Text has natural boundaries (characters, tokens or words), while user behavior is a continuous stream with irregular patterns across time. A user might be active daily, then disappear for months. We thus need to come up with a tokenisation scheme that can capture the user’s activity in a time window t.
2. Multi-Modality of tokens: Language models operate on discrete tokens from a fixed vocabulary. User behavior data consists of continuous numerical features (session lengths, activity counts) requiring different embedding strategies. Additionally user activity in a time window t is informative only in the event context of time t. For example, high activity on days where there are matches of a specific team implies that the user is a fan of that team and is likely to be engaged only when their team plays. This means that tokens need to be multi-modal, i.e. they need to contain information about the user’s activity in the time window t as well as the event context of time t.
4. Known Future Context: Unlike language generation, user behavior prediction often involves known future events (scheduled matches, planned promotions) that influence predictions. Decoder-only models like GPT are not able to attend to the future context since they are trained to predict the next token in a sequence given the previous tokens.
5. Multi-Task Objectives: Unlike text where Transformers are trained to predict the next token in a sequence - a single, multi-class classification problem, user behaviour prediction is not a single task but a multi-task problem. Typically we want to predict multiple tasks (churn, revenue, engagement) with varying nature of predictions (binary, continuous) for different time horizons from the same user representation.
@@ 3.1 Problem Formulation
  • $\mathbf{U}^{hist}\in\mathbb{R}^{T_{hist}\times d_{u}}$: Historical user transaction features over $T_{hist}$ days. For each day $t$, $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ contains the user’s aggregated transaction features for that day.
@@
The model outputs future behavior tokens $\hat{\mathbf{U}}^{fut}\in\mathbb{R}^{T_{fut}\times d_{u}}$ for each of the next $T_{fut}$ days. For each future day $t$, $\hat{\mathbf{u}}_{t}\in\mathbb{R}^{d_{u}}$ contains the predicted user’s aggregated transactions for that day. Each token contains the full set of predicted user behaviours, mirroring the dimensionality and structure of the input user features $\mathbf{U}^{hist}$.
@@ 3.2 Sequence Construction and Temporal Windows
We construct training sequences using fixed temporal windows of $T_{hist}=360$ days for historical context and $T_{fut}=7$ days for future predictions. The 360-day window captures the cyclical nature of the sports calendar, allowing the model to learn annual patterns such as recurring tournaments and league seasons, while the 7-day horizon aligns with practical short-term prediction needs in applications like churn prevention and product recommendations. User activity is inherently sparse, so for days without transactions we set $\mathbf{u}_{t}=\mathbf{0}$ while still providing event context $\mathbf{s}_{t}$ and static features $\mathbf{x}^{static}$, enabling the model to learn disengagement patterns as a function of available events. For users with fewer than 360 days on the platform, we pad the sequence with a learned padding token $\mathbf{u}_{t}=\mathbf{u}_{pad}$ (distinct from $\mathbf{0}$) for days before registration, allowing the model to distinguish between pre-registration periods and days with zero activity.
@@ 3.3 LUMOS Architecture Overview
LUMOS employs an encoder-decoder transformer architecture with specialized components for user behavior modeling. Figure 2 illustrates the complete architecture with three main stages: (1) token construction combining user transactions, event context, and static features into unified daily tokens, (2) the encoder processing historical token sequences through self-attention to produce contextualized representations, and (3) the decoder using cross-attention where future event context tokens query the encoded historical representations to generate future-conditioned embeddings, which are then mapped to predicted user behavior dimensions via output projection layers. We now detail each component.
Figure 2: LUMOS architecture overview.
@@ 3.3.1 Multi-Modal Token Construction
To tackle the first two challenges mentioned in the background section, namely user behaviour tokenisation and the multi-modality of tokens, we adopt a novel tokenization scheme. We aggregate the user activity $\mathbf{U}^{hist}$ and raw event context vectors $\mathbf{S}^{hist}$ at a daily level to yield daily tokens which are discrete in time and equally spaced. This addresses the first challenge by yielding discrete tokens.
The second challenge is to address the multi-modality of tokens. Since we have three modalities of data—namely user activity, event context, and static features $\mathbf{x}^{static}$—we embed each modality using a configurable MLP and concatenate the embeddings. Then we pass this concatenated embedding through a final MLP—called the cross-modal interaction layer—to project it to the model dimension $d_{model}$. Although user demographic features are invariant to time, we broadcast them across all time steps. Formally, for each timestep $t$, we embed the three modalities as $\mathbf{h}^{static}=\text{MLP}_{static}(\mathbf{x}^{static})$, $\mathbf{h}_{t}^{user}=\text{MLP}_{user}(\mathbf{u}_{t})$, and $\mathbf{h}_{t}^{event}=\text{MLP}_{event}(\mathbf{s}_{t})$, then construct the token as:
@@
Future Token Construction. Future tokens are constructed differently since we do not yet observe user behavior. We embed future event context via $\mathbf{z}_{t}^{fut}=\text{MLP}_{event}^{fut}(\mathbf{s}_{t}^{fut})$ using a separate embedding layer to learn distinct representations for past versus future event context.
@@ 3.3.2 Temporal Encoding
Temporal position is critical for user behavior modeling, as recent activity typically matters more than distant history, though specific past events (such as last year’s similar events) can be highly relevant. While language models conventionally rely on sinusoidal additive embeddings [15] or multiplicative rotation-based encodings like RoPE [12], the nature of position embeddings has been shown to be highly domain-specific, with time series models often benefiting from specialized temporal encodings [5]. Through rigorous experimentation, we found that learned position embeddings yield the best performance compared to sinusoidal embeddings, RoPE, or time series-specific encodings. The position embeddings are initialised to 0, are added to the token sequence and are learned during training.
@@ 3.3.5 User Behaviour Token Prediction
Each predicted token $\hat{\mathbf{u}}_{t}\in\mathbb{R}^{d_{u}}$ maintains identical dimensionality to the historical user tokens, containing the same aggregated transaction features across all behavioural dimensions.
@@ 3.5 Training Pipeline and Infrastructure
Training LUMOS at scale requires a sophisticated data pipeline and distributed training infrastructure to handle the massive volume of user behavioral data. Figure 3 illustrates the complete training pipeline showing the data flow from raw user transactions through PySpark preprocessing, Delta Lake storage, streaming data loading via Accio, and distributed training across multiple GPUs.
Figure 3: LUMOS training pipeline.
Data Scale and Preparation. Our training dataset comprises 250 million users over extended temporal windows, resulting in 275 billion user activity tokens. Each training observation constitutes a historical sequence of up to 360 daily tokens, with each token containing aggregated transaction features across multiple dimensions ($d_{u}=13$ behavioral dimensions). The complete dataset spans  50 terabytes when including historical user transactions, event context sequences, and static features. To handle this scale, we employ Apache Spark for distributed data preprocessing. The resulting feature-engineered dataset is stored in Delta Lake tables, which provide ACID transaction guarantees, schema evolution, and efficient time-travel capabilities for reproducible experiments.
@@
Training Duration and Compute. A complete training run of the full LUMOS model (with 6 encoder layers, 1 decoder layer, $d_{model}=512$, and 8 attention heads) requires approximately 72 hours (3 days) on our 8-GPU setup. This translates to 576 A100 GPU-hours per training run ($8\text{ GPUs}\times 72\text{ hours}$). We employ mixed-precision training using BF16 (Brain Floating Point) [3] to accelerate computation and reduce memory footprint while maintaining numerical stability.
@@ 4.1 Experimental Design
We designed our evaluation to assess LUMOS’s ability to predict multiple user behavior attributes simultaneously using a single unified model. The model we trained using LUMOS predicts 13 key user behavior attributes spanning engagement patterns, activity levels, and platform interactions on a daily aggregated basis. For 5 of these 13 attributes, we had established baseline models that enabled quantitative comparison: specialized deep neural networks with extensive hand-crafted features. We compare against these production baselines rather than sequential recommendation models (SASRec, BERT4Rec) due to fundamental architectural incompatibilities: those models lack cross-attention decoders for conditioning on known future events ($\mathbf{S}^{fut}$) and operate on discrete item vocabularies rather than continuous multi-modal daily representations. Adapting them would require such extensive modifications that they would no longer serve as representative baselines. We evaluated both offline model performance metrics (3 binary classification tasks using ROC-AUC, 2 continuous prediction tasks using MAPE) and online business impact through an A/B test.
The offline evaluation uses a large-scale dataset from a major sports engagement platform comprising 275 billion user activity tokens from 250 million users over four years (2021-2024). The data consists of user transaction sequences $\mathbf{U}^{hist}\in\mathbb{R}^{T_{hist}\times d_{u}}$, external event context $\mathbf{S}^{hist}\in\mathbb{R}^{T_{hist}\times d_{s}}$ (sports calendar), and static user attributes $\mathbf{x}^{static}\in\mathbb{R}^{d_{static}}$. Due to the large scale, we used only 1% of the data as holdout set for offline evaluation.
To validate whether offline performance improvements translate to measurable business impact, we deployed LUMOS predictions to feed into an intelligent offers engine designed to enhance user engagement through personalized offers. Previously, the offers engine relied on a combination of legacy specialized baseline models. We designed a controlled A/B test comparing offers engine performance using LUMOS predictions against using the legacy specialized models. For the experiment, we created two random, mutually exclusive cohorts of approximately 700,000 users each: (1) Target group receiving offers based on LUMOS predictions, and (2) Control group receiving offers based on legacy specialized model predictions. The experiment ran for 18 days. The primary metrics measured were Daily Active Users (DAU) and cost efficiency of the offers strategy.

### v2

@@ Abstract
User behavior prediction at scale remains a critical challenge for online B2C platforms. Traditional approaches rely heavily on task specific models and domain specific feature engineering. This is time-consuming, computationally expensive, and requires domain expertise and therefore not scalable. We present LUMOS (Large User MOdel Series), a transformer-based architecture that eliminates task specific models and manual feature engineering by learning multiple tasks jointly using only raw user activity data. LUMOS introduces a novel cross-attention mechanism that conditions predictions on future known events (e.g., holidays, sales, etc.), enabling the model to predict complex behavior patterns like ”how will upcoming holidays affect user engagement?” The architecture also employs multi-modal tokenization, combining user activities, event context, and static user demographic attributes into rich representations processed through specialized embedding pathways.
Through extensive experiments on a production dataset spanning 1.7 trillion user activity tokens from 250 million users, we demonstrate that LUMOS achieves superior performance compared to traditional task-specific models. Across 5 tasks with established baselines, we achieve an average improvement of 0.025 in ROC-AUC for binary classification tasks and 4.6% reduction in MAPE for regression tasks. Online A/B testing validates these improvements translate to measurable business impact with a 3.15% increase in Daily Active Users.
@@ 1 Introduction
We propose a shift in user behavior modelling inspired by the trends in NLP where the field moved from task-specific models and feature based approaches(like bag of words and TF-IDF) to foundation models like BERT [1] and GPT [5] which are not just easier to train(in terms of pre-processing steps) but also achieve better performance when data scale is sufficient.
Figure 1: Paradigm shift from multiple specialized models to a single unified model. Traditional approaches require separate models for each prediction task (left), while LUMOS employs a single powerful model capable of handling multiple tasks simultaneously (right).
In this spirit, we present LUMOS, a transformer-based architecture that can be trained to predict future user behaviour using raw sequence of past behaviour. It eliminates manual feature engineering by learning end-to-end representations from raw user activity data. Instead of predicting a single user behaviour attribute for a pre-specified time horizon, LUMOS predicts a sequence of user behaviour tokens for multiple future time steps. Each predicted token contains the full set of predicted user behaviours precluding the need for multiple specialized models. Experiments on our production dataset show that LUMOS achieves an average improvement of 0.025 in AUC for classification tasks and 4.6% reduction in MAPE for regression tasks compared to the baseline task-specific models.
LUMOS uses a cross-attention mechanism between historical user behavior and future event context to answer nuanced questions like ”Given that the IPL final is next Sunday, which days in the user’s history are most relevant for predicting their engagement?” Our experiments show that the model correctly attends not only to recent activity but also to specific historical events like previous year’s IPL finals or World Cup matches, learning these patterns entirely from data without explicit programming. This is crucial for event driven platforms like sports streaming platforms or ticketing services.
@@ 2 Background
User behavior modeling encompasses a range of tasks critical to online platforms: predicting churn, forecasting engagement levels, anticipating user actions, and understanding activity patterns [7]. These predictions typically span multiple time horizons (7-day, 28-day, annual) and serve different business functions from personalized marketing to user retention strategies.
@@
While transformers excel at natural language processing, adapting them to user behavior prediction presents unique challenges:
1. Continuous vs. Discrete Temporal Patterns: Text has natural boundaries (characters, tokens or words), while user behavior is a continuous stream with irregular patterns across time. A user might be active daily, then disappear for months. We thus need to come up with a tokenization scheme that can capture the user’s activity in a time window t.
2. Multi-Modality of tokens: Language models operate on discrete tokens from a fixed vocabulary. User behavior data consists of continuous numerical features (session lengths, activity counts) requiring different embedding strategies. Additionally user activity in a time window t is informative only in the event context of time t. For example, high activity on a days where there are matches of a specific team implies that the user is a fan of that team and is likely to be engaged only when their team plays. This means that tokens need to be multi-modal, i.e. they need to contain information about the user’s activity in the time window t as well as the event context of time t.
4. Known Future Context: Unlike language generation, user behavior prediction often involves known future events (scheduled matches, planned promotions) that influence predictions. Decoder only models like GPT are not able to attend to the future context since they are trained to predict the next token in a sequence given the previous tokens.
5. Multi-Task Objectives: Unlike text where Transformers are trained to predict the next token in a sequence - a single ,multi class classification problem, user behaviour prediction is not a single task but a multi-task problem. Typically we want to predict multiple tasks (churn, activity levels, engagement) with varying nature of predictions (binary, continuous) for different time horizons from the same user representation.
@@ 3.1 Problem Formulation
  • $\mathbf{U}^{hist}\in\mathbb{R}^{T_{hist}\times d_{u}}$: Historical user activity features over $T_{hist}$ days. For each day $t$, $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ contains the user’s aggregated activity features for that day.
@@
The model outputs future behavior tokens $\hat{\mathbf{U}}^{fut}\in\mathbb{R}^{T_{fut}\times d_{u}}$ for each of the next $T_{fut}$ days. For each future day $t$, $\hat{\mathbf{u}}_{t}\in\mathbb{R}^{d_{u}}$ contains the predicted user’s aggregated activities for that day. Each token contains the full set of predicted user behaviours, mirroring the dimensionality and structure of the input user features $\mathbf{U}^{hist}$.
@@ 3.2 Sequence Construction and Temporal Windows
We construct training sequences using fixed temporal windows of $T_{hist}=360$ days for historical context and $T_{fut}=7$ days for future predictions. The 360-day window captures the cyclical nature of the sports calendar, allowing the model to learn annual patterns such as recurring tournaments and league seasons, while the 7-day horizon aligns with practical short-term prediction needs in applications like churn prevention and product recommendations. User activity is inherently sparse, so for days without activity we set $\mathbf{u}_{t}=\mathbf{0}$ while still providing event context $\mathbf{s}_{t}$ and static features $\mathbf{x}^{static}$, enabling the model to learn disengagement patterns as a function of available events. For users with fewer than 360 days on the platform, we pad the sequence with a learned padding token $\mathbf{u}_{t}=\mathbf{u}_{pad}$ (distinct from $\mathbf{0}$) for days before registration, allowing the model to distinguish between pre-registration periods and days with zero activity.
@@ 3.3 LUMOS Architecture Overview
LUMOS employs an encoder-decoder transformer architecture with specialized components for user behavior modeling. Figure 2 illustrates the complete architecture. We now detail each component.
Figure 2: LUMOS architecture showing the complete model pipeline from token construction to prediction. The token construction process (bottom) combines user activities, event context, and static features into unified daily tokens. The encoder processes the historical token sequence through self-attention, producing contextualized representations of the same dimension as the input tokens. The decoder uses cross-attention where future event context tokens query the encoded historical representations to generate future-conditioned embeddings. Finally, an output projection layer maps each decoder output to predicted user behavior dimensions $\mathbf{U}^{fut}$.
@@ 3.3.1 Multi-Modal Token Construction
To address the challenges of tokenizing continuous user behavior (Section 2), we adopt a novel multi-modal tokenization scheme. We aggregate user activity $\mathbf{U}^{hist}$ and event context $\mathbf{S}^{hist}$ at a daily level, yielding discrete, equally-spaced temporal tokens. Since we have three modalities—user activity, event context, and static features—we embed each modality separately using configurable MLPs, then fuse them via a cross-modal interaction layer.Each modality is processed by a dedicated MLP with configurable depth (1-3 layers, we use depth=2). These embeddings transform raw numerical features into learned representations:
$\mathbf{h}^{static}$ $=\text{MLP}_{static}(\mathbf{x}^{static})\in\mathbb{R}^{d_{static\_embed}}$ (1)
$\mathbf{h}_{t}^{user}$ $=\text{MLP}_{user}(\mathbf{u}_{t})\in\mathbb{R}^{d_{user\_embed}}$ (2)
$\mathbf{h}_{t}^{event}$ $=\text{MLP}_{event}(\mathbf{s}_{t})\in\mathbb{R}^{d_{supply\_embed}}$ (3)
Cross-Modal Interaction Layer. Unlike simple concatenation, we pass the concatenated embeddings through a final projection MLP that learns interactions between modalities. For example, a user’s high activity level during IPL finals has different meaning than the same activity during off-season. This layer captures such contextual dependencies:
@@
where $\cdot$$\cdot$] denotes concatenation and static features are broadcast across all timesteps.
Future Token Construction. For future timesteps, we only have event context: $\mathbf{z}_{t}^{fut}=\text{MLP}_{event}^{fut}(\mathbf{s}_{t}^{fut})$, using a separate embedding layer to learn distinct representations for future versus historical event context. The token construction pipeline is illustrated in Figure 2.
@@ 3.3.2 Temporal Encoding
Temporal position is critical for user behavior modeling, as recent activity typically matters more than distant history, though specific past events (e.g., last year’s tournament finals) can be highly relevant. While language models conventionally rely on sinusoidal embeddings [11] or rotation-based encodings like RoPE [9], position encoding behavior is highly domain-specific.
Through systematic comparison (Table 3 in Appendix 0.A), we found learned position embeddings yield the best performance. The embeddings are initialized to zero, added to token sequences, and optimized during training. Fourier analysis of the learned embeddings reveals they primarily encode low-frequency temporal patterns (see Appendix 0.A for detailed frequency analysis). This suggests that for user behavior, broad temporal context (recent week vs. distant past) matters more than precise day-level positioning, contrasting with language models where exact token position is critical for syntax and semantics.
@@ 3.3.5 User Behaviour Token Prediction
Each predicted token $\hat{\mathbf{u}}_{t}\in\mathbb{R}^{d_{u}}$ maintains identical dimensionality to the historical user tokens, containing the same aggregated activity features across all behavioural dimensions.
@@ 3.4 Training Objective
3.5 Model Specifications
Table 1 provides complete specifications for the production LUMOS model.
Table 1: LUMOS Model Card
Architecture
Type Encoder-Decoder Transformer with Cross-Attention
Total Parameters 7.23M
Encoder Layers 6 layers, 8 heads, $d_{model}=512$, $d_{ff}=2048$
Decoder Layers 6 layers, 8 heads (cross-attention only)
Positional Encoding Learned embeddings
Activation SwiGLU (transformer), ReLU (projection heads)
Normalization Pre-LayerNorm
Input/Output Specifications
Historical Window $T_{hist}=360$ days
Future Window $T_{fut}=7$ days
User Features $d_{u}=13$ dimensions
Event Context Features $d_{s}=32$ dimensions
Static Features $d_{static}=9$ dimensions
Token Embedding Dims user=128, event context=64, static=32
Output Predictions $(T_{fut},d_{u})$ - 3 binary, 10 continuous
Training Configuration
[…]
Batch Size 512 per GPU $\times$ 8 GPUs = 4096 effective
Precision BF16 mixed precision
[…]
Training Time 72 hours (576 A100 GPU-hours)
Dataset 250M users, $\sim$48TB processed
Auxiliary Outputs. Beyond behavioral predictions, LUMOS generates reusable embeddings: (1) dynamic user embeddings ($d_{model}=512$) from encoder outputs, (2) static user embeddings ($d_{static\_embed}=32$) for cold-start scenarios, and (3) event context embeddings ($d_{supply\_embed}=64$) representing daily event context. See Appendix 0.B for extraction methods and evaluation.
@@ 3.6 Training Pipeline and Infrastructure
Training LUMOS at scale requires a sophisticated data pipeline and distributed training infrastructure to handle the massive volume of user behavioral data. Figure 3 illustrates the complete training pipeline from raw data to model deployment.
Figure 3: LUMOS training pipeline showing the data flow from raw user activities through PySpark preprocessing, Delta Lake storage, streaming data loading via Accio, and distributed training across multiple GPUs.
Data Scale and Preparation. Our training dataset comprises 250 million users over extended temporal windows, resulting in  1.7 trillion user activity tokens. Each training observation constitutes a historical sequence of up to 360 daily tokens, with each token containing aggregated activity features across multiple dimensions ($d_{u}=13$ behavioral dimensions). The complete dataset spans  50 terabytes when including historical user activities, event context sequences, and static features. To handle this scale, we employ Apache Spark for distributed data preprocessing. The resulting feature-engineered dataset is stored in Delta Lake tables, which provide ACID transaction guarantees, schema evolution, and efficient time-travel capabilities for reproducible experiments.
@@
Training Duration and Compute. A complete training run of the full LUMOS model (with 6 encoder layers, 6 decoder layers, $d_{model}=512$, and 8 attention heads) requires approximately 72 hours (3 days) on our 8-GPU setup. This translates to 576 A100 GPU-hours per training run ($8\text{ GPUs}\times 72\text{ hours}$). We employ mixed-precision training using BF16 (Brain Floating Point) [3] to accelerate computation and reduce memory footprint while maintaining numerical stability.
@@ 4.1 Experimental Design
We designed our evaluation to assess LUMOS’s ability to predict multiple user behavior attributes simultaneously using a single unified model. THe model we trained using LUMOS predicts 13 key user behavior attributes spanning engagement patterns, activity levels, and platform interactions on a daily aggregated basis. For 5 of these 13 attributes, we had established baseline models that enabled quantitative comparison: specialized Random Forest ensembles and deep neural networks with extensive hand-crafted features. We evaluated both offline model performance metrics (3 binary classification tasks using ROC-AUC, 2 continuous prediction tasks using MAPE) and online business impact through an A/B test.
The offline evaluation uses a large-scale dataset from Dream11 comprising 1.7 trillion user activity tokens from 250 million users over four years (2021-2024). The data consists of user activity sequences $\mathbf{U}^{hist}\in\mathbb{R}^{T_{hist}\times d_{u}}$, external event context $\mathbf{S}^{hist}\in\mathbb{R}^{T_{hist}\times d_{s}}$ (sports calendar), and static user attributes $\mathbf{x}^{static}\in\mathbb{R}^{d_{static}}$. Due to the large scale, we used only 1% of the data as holdout set for offline evaluation.
To validate whether offline performance improvements translate to measurable business impact, we deployed LUMOS predictions to feed into Pacman—our proprietary intelligent offers engine designed to enhance user engagement through personalized offers. Previously, Pacman relied on a combination of legacy specialized baseline models. We designed a controlled A/B test comparing Pacman performance using LUMOS predictions against using the legacy specialized models. For the experiment, we created two random, mutually exclusive cohorts of approximately 700,000 users each: (1) Target group receiving offers from Pacman based on LUMOS predictions, and (2) Control group receiving offers based on legacy specialized model predictions. The experiment ran for 18 days. The primary metrics measured were Daily Active Users (DAU) and cost efficiency of the offers strategy.